OpenStudy (anonymous):

Help pwease I'm lostttttttt f(x) = x^2 + 4x - 21 = 24 the answer is (x-5)(x+9) .... Why can't it be (x+9)(x-5)?

OpenStudy (agent0smith):

(x-5)(x+9) and (x+9)(x-5) are the exact same expression, just written in a different order. You can change the order of multiplication... 3*5 is the same as 5*3, they're both 15.

OpenStudy (nincompoop):

it cane be, how do you know? do the FOIL method and expand

OpenStudy (anonymous):

sorry for the late response i keep getting disconnected from openstudy....

OpenStudy (anonymous):

so you're saying it work either way?

OpenStudy (anonymous):

nvm it doesn't work either way it has to be -9

OpenStudy (agent0smith):

Order does not matter in multiplication. Is 4*6 any different than 6*4?

OpenStudy (anonymous):

oh wait !!!!!!!!!!! i got it so it doesn't matter !!!! i just figured it out!!!

OpenStudy (anonymous):

x + 9 = 0 so its -9 and x - 5 = 0 so its -5!!!

OpenStudy (anonymous):

got it niceeee thanks !!!

OpenStudy (agent0smith):

x-5 = 0 means x = 5, not -5